Solve the separable differential equation


10x-8ysqrt(x^2 +1) * dy/dx =0




Subject to the initial condition: .y(0)=9



y=??

1
Expert's answer
2022-03-10T18:06:30-0500

Let us solve the differential equation

"10x-8y\sqrt{x^2 +1}\frac{ dy}{dx }=0,"

which is equivalent to

"\frac{10x dx}{ \sqrt{x^2 +1}}=8ydy."

It follows that

"\int\frac{10x dx}{ \sqrt{x^2 +1}}=\int8ydy."

Therefore, the general solution is

"10\sqrt{x^2+1}=4y^2+C."

Since "y(0)=9," we get "10=4\cdot 81+C."

Therefore, "C=-314."

Consequently, the particular solution is

"10\sqrt{x^2+1}=4y^2-314."